an alternating voltage ( in Volts) given by V= 200 sin (100t) is connected to 1F capacitor through an ideal ac ammeter in series. What shall be the reading of the ammeter and the average power consumed in the circuit .

Grade:

1 Answers

NISHARG GOLASH
10 years ago

Hi Amit

as 

V= 200 sin (100t)

So, V0 = 200 and frequency(w) = 100

Ammeter current = Vo/R

so, now we have to calculate R i.e capacitive resistance = 1/wc

R = 1/(100*1) 

R = 0.01 ohm

So, Ammeter reading = 200/0.01

I = 2000 A

And Avg Power = 1/2 * CV^2

 = 0.5* 1* 200^2

 = 20000 watt

 

please refer to theory on http://www.electronics-tutorials.ws/capacitor/cap_8.html for better understanding